# Integrating Large Language Models into a Tri-Modal Architecture for Automated Depression Classification

### v3

@@ II-A Mel Frequency Cepstral Coefficients in Audio-Based Models
Derived using the Mel Scale [29], MFCCs provide a compact representation of the human perception of sound. This allows for the low-level representation of timbre, pitch, and rhythm. Studies indicate such values are different between depressed and non-depressed patients [30]. (Wang et al.) [31] and (Taguchi et al.) [32] found certain MFCCs follow a consistent pattern that is only present in depressed patients.
@@ III-A DAIC-WOZ
Access to DAIC-WOZ was granted after signing an End User License Agreement (EULA) [^1] Zipped DAIC-WOZ files can be found after signing the necessary EULA at https://dcapswoz.ict.usc.edu/.
@@ 0 Therapist So how are you doing today?
The DAIC-WOZ contains transcripts as tab-separated .CSV files with the following columns: speaker, start_time, stop_time, and value. The provided utterances contained numerous errors that needed to be fixed. [^2] Documentation detailing how the original DAIC-WOZ transcriptions were created can be found at Preprocessing was applied using RegEx and NLP techniques to make text suitable to pass into the LLM. Refer to Table II for a sample of a transcript before pre-processing is applied.
@@ III-D1 Audio Segmentation
  • Pitch Shifting: Pitch was adjusted by N.N, N.N, and N.N in semitones.
@@ III-E1 Video Segmentation
In the proposed architecture, we analyze patient visual data only while they are speaking, not in all parts of the interview. Using the timestamps provided in the transcript, FAU data was cropped to exclude and discard segments where the therapist is speaking, or where nobody is speaking at all. The data was then sliced into 8-second segments in such a way that it aligned with the segmented audio data. Any remeaining segment under 8 seconds was discarded. After video segmentation was applied, there were an equal number of video and audio samples for each interview.
@@ V Results
Predicted Actual D ND total D 13 (27.66) 1 (2.13) 92.85 ND 6 (12.77) 27 (57.45) 81.81 total 68.41 96.42 Predicted Actual D ND total D 52 (27.51) 13 (6.88) 79.99 ND 4 (2.12) 120 (63.49) 96.77 total 92.85 90.22

### v4

@@ II-A Mel Frequency Cepstral Coefficients in Audio-Based Models
Derived using the Mel Scale [29], MFCCs provide a compact representation of the human perception of sound. This allows for the low-level analysis of timbre, pitch, and rhythm. Studies indicate such values are different between depressed and non-depressed patients [30]. (Wang et al.) [31] and (Taguchi et al.) [32] found certain MFCCs follow a consistent pattern that is only present in depressed patients.
@@ III-A DAIC-WOZ
Access to DAIC-WOZ was granted after signing an End User License Agreement (EULA) [^1] Zipped DAIC-WOZ files can be found after signing the necessary EULA at https://dcapswoz.ict.usc.edu/..
@@ 0 Therapist So how are you doing today?
The DAIC-WOZ contains transcripts as tab-separated .CSV files with the following columns: speaker, start_time, stop_time, and value. The provided utterances contained numerous errors that needed to be fixed. [^2] Documentation detailing how the DAIC-WOZ can be found online. Preprocessing was applied using RegEx and NLP techniques to make text suitable to pass into the LLM. Refer to Table II for a sample of a transcript before pre-processing is applied.
@@ III-D1 Audio Segmentation
  • Pitch Shifting: Pitch was adjusted by 0.5, 2.0, and 2.5 in semitones.
@@ III-E1 Video Segmentation
In the proposed architecture, we analyze patient visual data only while they are speaking, not in all parts of the interview. Using the timestamps provided in the transcript, FAU data was cropped to exclude and discard segments where the therapist is speaking, or where nobody is speaking at all. The data was then sliced into 8-second segments in such a way that it aligned with the segmented audio data. Any remaining segment under 8 seconds was discarded. After video segmentation was applied, there were an equal number of video and audio samples for each interview.
@@ V Results
Predicted Actual D ND D 16 (34.04) 1 (2.13) ND 3 (6.38) 27 (57.45) Predicted Actual D ND total D 52 (27.51) 13 (6.88) 79.99 ND 4 (2.12) 120 (63.49) 96.77 total 92.85 90.22